Determining Amounts to be Paid on Invoices Determine the amount to be paid in full settlement of each of the following invoices, assuming that credit for returns and allowances was received prior to payment and that all invoices were paid within the discount period: Merchandise Freight Paid by Seller Customer Returns and Allowances a. $13,400 - FOB destination, n/30 $700 b. 9,500 $400 FOB shipping point, 2/10, n/30 1,100 c. 7,800 - FOB shipping point, 2/10, n/30 700 d. 3,700 100 FOB shipping point, 1/10, n/30 500 e. 1,200 - FOB destination, 2/10, n/30 - a. $ 12,700 b. $ 8,400 c. $ 7,100 d. $ 3,200 e. $ 1,176
Answer #1

Calculate amount paid

a) 13400-700 = 12700
b) (9500-1100)*98%+400 = 8632
c) (7800-700)*98% = 6958
d) (3700-500)*99%+100 = 3268
e) 1200*98% = 1176